# https://yh2371.github.io/como

### v1

@@ Abstract
Text-to-motion models excel at efficient human motion generation, but existing approaches lack fine-grained controllability over the generation process. Consequently, modifying subtle postures within a motion or inserting new actions at specific moments remains a challenge, limiting the applicability of these methods in diverse scenarios. In light of these challenges, we introduce CoMo, a Controllable Motion generation model, adept at accurately generating and editing motions by leveraging the knowledge priors of large language models (LLMs). Specifically, CoMo decomposes motions into discrete and semantically meaningful pose codes, with each code encapsulating the semantics of a body part, representing elementary information such as “left knee slightly bent”. Given textual inputs, CoMo autoregressively generates sequences of pose codes, which are then decoded into 3D motions. Leveraging pose codes as interpretable representations, an LLM can directly intervene in motion editing by adjusting the pose codes according to editing instructions. Experiments demonstrate that CoMo achieves competitive performance in motion generation compared to state-of-the-art models while, in human studies, CoMo substantially surpasses previous work in motion editing abilities.
@@ 1 Introduction
To address these challenges, we propose the Controllable Motion Generation model (CoMo). As shown in Figure 1, CoMo can generate high-quality human motions from a broad range of text (e.g., “A person throws a punch.”). Users can control the generation by altering the descriptions for each body part, e.g., switching the left and right arm descriptions changes the punching hand. More importantly, CoMo enables detailed editing across frames and joints (e.g., Hunch over more.), adding actions (e.g., Squat at the end.), and varying speed and even emotion (e.g., more dramatic), as depicted in Figure 6.
CoMo achieves these capabilities by representing motions as interpretable “pose codes”, with each code defining the state of a specific body part at a given moment, e.g., right arm straight. As demonstrated in Figure 2, our method starts by factorizing a motion sequence into a series of temporal states, each composed of pose codes that describe the motion’s kinematic characteristics. This structure allows for precise encoding of human motion sequences in time and fine-grained control of kinematic joint states. Since these encodings are explicitly interpretable, people can interact with and modify the sequences intuitively. We show that this process can also be made instructable with natural language by allowing large language models (LLM) to edit the codes as well.
Leveraging pose codes as interpretable motion representations, the three main components of CoMo work jointly to effectively generate and edit motion: (1) The Motion Encoder-Decoder parses motions into sequences of pose codes and reconstructs them back into motions by learning a decoder; (2) The Motion Generator, a transformer-based model, generates pose codes conditioned on text inputs and LLM-generated fine-grained descriptions; (3) The Motion Editor uses LLMs to modify and refine the pose codes based on editing instructions. CoMo allows for intuitive, language-controlled adjustments to the motion sequences, both temporally and kinematically, closely aligning the generated motions with users’ creative intentions and the nuances expressed in their textual descriptions, making the process user-friendly and adaptable to diverse applications.
@@ 2 Related Work
Fine-grained motion generation has gained significant interest due to its extensive practical applications. However, current methods have not fully integrated spatial and temporal details. TEACH [1] demonstrates improvements in motion smoothness by incorporating extensive temporal annotations but requiring additional labeled data and overlooks lower-level spatial details. [17, 30, 2] leverage LLMs to generate detailed text descriptions for whole-body motions and individual body parts, aiming to map text to spatiotemporal details. However, without explicit supervision, these fine-grained details may not align with the corresponding motion. GraphMotion [16] proposes a hierarchical semantic graph that enforces coarse-to-fine topology within the text-to-motion diffusion process. However, constructing the hierarchical graph depends on the semantic parsing of text details and thus may struggle for ambiguous inputs. Also, the graph nodes do not model the lower-level spatial characteristics and relations. To this end, we propose body-part-specific semantic pose codes for achieving fine-grained representation of motion sequences. In addition, these pose codes are used as context to guide LLMs in generating motion-coherent descriptions for different body parts to enhance the connection between text and fine-grained motion details.
Motion Editing enables users to interactively refine generated motions to suit their expectations. PoseFix [6] automates 3D pose and text modifier generation for supervised editing, but frame-by-frame pose edits lack efficiency and temporal consistency. TLControl [35] allows motion editing using joint-level trajectories but relies on high-quality trajectory inputs and is not intuitive for interactive motion editing. [13] employs an autoencoder to optimize a motion manifold under positional, bone length, and trajectory constraints for smooth edits. Diffusion-based approaches [18, 9, 37] can achieve zero-shot spatiotemporal editing by infilling particular joints or frames, which may form unnatural discontinuities. Recently, FineMoGen [43] optimizes global attention for fine-grained editing but generates new sequences for each edit, limiting motion consistency. In our approach, we encode motions into semantic pose codes, which serve as context to prompt an LLM to edit the original motion directly, ensuring motion consistency during the process.
@@ 3 Method
Figure 2: Overview of CoMo for text-driven motion generation. Motion Encoder-Decoder (left) utilizes a predefined codebook to encode motions into pose codes and learns a decoder to reconstruct the motions. Motion Generator (right), a transformer-based model, predicts pose codes autoregressively, conditioned on the text descriptions and LLM-generated fine-grained keywords. The generated pose codes are then decoded back into motions.
@@ 3.1 Motion Encoder-Decoder
As shown in the left panel of Figure 2, given $T$ frames of motion $X=\left\{x_{i}\right\}_{i=1}^{T}$, the motion encoder $\mathcal{E}$ projects them into a sequence of pose codes using a codebook, denoted as $Z=\mathcal{E}\left(X\right)$, where $Z=\left\{z_{i}\right\}_{i=1}^{L}$. Here, $L=T/l$ represents the length of the pose code sequence, and $l$ is the downsampling rate. The decoder $\mathcal{D}$ then decodes the pose codes back into motions, expressed as $X_{\text{rec}}=\mathcal{D}(\hat{Z})$, where $\hat{Z}$ stands for the latent features of pose codes. In the following, we will explain how to build the pose codebook, encode the motions using these codes, and learn the decoder to recover the motions.
Pose Codebook. Unlike [40], which employs an autoencoder to obtain implicit motion representations, CoMo predefines a pose codebook with semantic elements to achieve interpretable motion factorization across time and space. Following PoseScripts [5], we construct a pose codebook with $N$ codes, $\mathcal{C}=\left\{c_{n}\right\}_{n=1}^{N}$, with $c_{n}\in\mathbb{R}^{d_{c}}$, where $d_{c}$ is the dimension of the learnable codebook entry. Each code $c_{n}$ is associated with a semantic meaning, representing a state of a body part or the spatial relationships between body parts, e.g., “left knee slightly bent”, “left hand and left foot close”. These codes are further grouped into $K$ pose categories, each encompassing different states of the same body parts, e.g., the pose category “left knee angle” includes codes like “left knee slightly bent”, “left knee partially bent”, etc.
Motion Encoder. Given the codebook $\mathcal{C}$ with $N$ pose codes across $K$ categories, we encode motions into $K$-hot $N$ dimensional vectors, denoted as $Z=\mathcal{E}\left(X\right)=\left\{z_{i}\right\}_{i=1}^{L}$, where $z_{i}\in\mathbb{R}^{N}$. At each time step $i$, only one pose code per pose category is activated (set to $1$ in the vector), i.e., $K$ elements of $z_{i}$ are set to $1$, while the others are 00.
To determine whether a pose code $c\in\mathcal{C}$ applies to a motion $x$, we use an off-the-shelf skeleton parser [5], denoted as $\mathcal{P}(c,x)\rightarrow\{0,1\}$. The parser $\mathcal{P}$ analyzes the 3D joint positions of a skeleton in SMPL format [20], evaluating whether the pose meets specific heuristic threshold conditions for a given pose code. For instance, if the angle formed by the left shoulder, elbow, and wrist joints is less than 20 degrees, the code “left arm completely bent” is true. Therefore, instead of training an encoder model, we can explicitly factorize the motion sequence as:
@@
We define $V(X)=\left\{x_{i+1}-x_{i}\right\}_{i=1}^{T-1}$ as the velocity of the $T$-frame motion sequence $X$. The motion reconstruction objective is formulated using the smooth L1 loss $\mathcal{L}_{1}$:
@@
where the hyperparameter $\lambda$ [^1] Following [40], we set $\lambda$ to 0.1. balances the velocity and motion loss.
@@ 3.2 Motion Generator
As illustrated in the right panel of Figure 2, the Motion Generator, conditioned on text input, aims to generate a sequence of pose codes, which will be decoded into motions.
Utilizing the pose codebook, we map a sequence of motions $X=\{x_{i}\}_{i=1}^{T}$ to a sequence of $K$-hot, $N$-dimensional vectors $Z^{1:L}=\left\{\{z_{i}^{n}\}_{n=1}^{N}\right\}_{i=1}^{L}$, where $z_{i}^{n}$ is an indicator function that is activated if the corresponding pose code $c_{n}$ is true at the time index $i$. To denote the end of a motion sequence, we append an <End> code to each latent vector, activated when the motion is stopped, and the dimension of each latent vector $z_{i}$ becomes $N+1$.
@@
To achieve this goal, we adopt a decoder-only transformer architecture with causal self-attention [40]. [^2] Following the method for processing image patches in Vision Transformers [7], a linear projection layer is adopted to project the $K$-hot vectors before input into the transformer architecture. The likelihood of the full sequence is
@@
The predicted sequences of $K$-hot vectors are mapped back to their corresponding pose codes, which can then be decoded to a motion sequence through the decoder $\mathcal{D}$.
Fine-grained Keywords. To help the model capture more fine-grained details in motion generation, we enhance the text description $t$ by using GPT-4 [24] to generate a keyword for each of the 10 body parts [^3] The 10 body parts are head, torso, left arm, right arm, left hand, right hand, left leg, right leg, left feet, right feet. and a keyword to describe the overall mood of the motion. For example, as shown in Figure 2, GPT-4 generates “bend at the knee”, “torso leaning forward”, etc., as the keywords for the motion crouching. We use CLIP [27] to extract the text embeddings of the keywords and original description, which serve as the conditions for the Motion Generator. The embedding of the original text, followed by the embeddings of 11 keywords, forms the initial token sequence that conditions the subsequent motion generation. The effectiveness of incorporating these keywords to improve generation performance is validated in Table 3.
@@ 3.3 Motion Editor
The edited segments of the pose codes are seamlessly integrated with their unedited counterparts, forming the complete edited sequence of pose codes. After the motion editing process, this revised pose code sequence is passed through the decoder $\mathcal{D}$ to reconstruct the final edited motion sequence.
@@ 4 Evaluation
In this section, we evaluate CoMo from two perspectives: (1) experiments to demonstrate that CoMo achieves state-of-the-art performance on text-driven motion generation (Sec 4.1), and (2) human evaluation to showcase that CoMo is superior over existing methods in motion editing (Sec 4.2). We also conduct comprehensive ablation studies (Sec 4.3) to validate the effectiveness of our model design.
@@ 4.1 Experiments on Motion Generation
Method R-Precision normal-$\uparrow$ FID normal-$\downarrow$ MM-DIST normal-$\downarrow$ Diversity normal-$\uparrow$ MModality normal-$\uparrow$
@@
Method R-Precision normal-$\uparrow$ FID normal-$\downarrow$ MM-DIST normal-$\downarrow$ Diversity normal-$\uparrow$ MModality normal-$\uparrow$
@@
Table 2: Comparison with the state-of-the-art methods on KIT [26] test set. The best performance is bold, and the second best is underlined, the third best is itallic.
CoMo achieves competitive results on text-driven motion generation. Table 1 and Table 2 present the results of text-driven motion generation on the HumanML3D and KIT-ML datasets, respectively. CoMo attains either the best or the second-best performance on HumanML3D across five metrics and ranks within the top three in six metrics on KIT. Compared to existing state-of-the-art methods, CoMo not only achieves competitive motion fidelity (FID) but also enhances motion diversity and consistency between generated motion and text descriptions (R-Precision, MM-DIST, Diversity). MModality measures the diversity of motions generated from the same text description. Due to the use of semantically meaningful pose codes, CoMo builds a stronger binding between the text description and the corresponding generated motion sequence, which potentially causes a slightly lower MModality score compared to prior methods as a trade-off for consistency. More importantly, our approach provides an intuitive interface for LLM-based zero-shot motion editing, which will be evaluated in Section 4.2 through a user study.
@@ 4.2 Human Evaluation on Motion Editing
Baselines. To form strong baselines, we prompt GPT-4 to generate an updated motion description in the style of HumanML3D annotations using the original description and edit instructions as context [^7] Prompts and updated descriptions are available in the Appendix.. Using the updated description, we then use T2M-GPT and FineMoGen to generate the edited motion sequence. For CoMo, we decompose the source motion sequence as pose codes and prompt GPT-4 to edit the code sequence based on the original description, edit instruction and pose code semantics without fine-tuning. The edited code sequence is decoded to reconstruct the edited motions.
@@
Humans prefer CoMo for Motion Editing. We present the average percentage of users’ preferences for CoMo versus T2M-GPT and FineMoGen in Figure 5. The results show a clear preference (over 70% on average) for motion editing with CoMo, especially in scenarios that involve modifying fine-grained motion details (such as body part modification and add/delete action). As depicted in Figure 6, while T2M-GPT and FineMoGen can produce motion sequences relevant to the updated description, they often struggle to generate well-aligned details from scratch without any spatial-temporal context of the source motion. In contrast, CoMo leverages the semantics of pose codes to interpret the source motion effectively and achieve detailed motion editing. The observed decrease in preference for holistic motion edits, which involve overall changes in emotion or speed, suggests that textual descriptions may provide more guidance in creating comprehensive edits that impact multiple aspects of a motion sequence. With its capability for high-level text-guided generation and iterative fine-grained pose code editing, CoMo emerges as a competitive approach for text-driven, controllable motion generation.
Figure 6: Qualitative examples of Motion Editing on the HumanML3D test set [10]. The green words and boxes highlight successful motion edits. The red words and boxes identify misalignments between the edited and source motions. Compared to other state-of-the-art methods, our approach can achieve accurate motion edits while preserving key characteristics of the source motion.
4.3 Ablation
@@
Top-1 normal-$\uparrow$ FID normal-$\downarrow$ MM-DIST normal-$\downarrow$ Diversity normal-$\uparrow$ Top-1 normal-$\uparrow$ FID normal-$\downarrow$ MM-DIST normal-$\downarrow$ Diversity normal-$\uparrow$
@@
Codebook Size. The number of pose codes is correlated to how fine-grained the heuristic thresholds of the skeleton parser are. More fine-grained thresholds capture more details but come at the expense of computational speed and learning complexity. We evaluate the motion reconstruction quality for different codebook sizes by varying the angle and distance threshold settings, with finer thresholds corresponding to larger codebooks. Results in Table 4 indicate that our current codebook setting of $392$ pose codes best balances complexity and reconstruction quality.
@@
Top-1 normal-$\uparrow$ FID normal-$\downarrow$ MM-DIST normal-$\downarrow$ Diversity normal-$\uparrow$
@@
Sampling Rate ($l$) Top-1 normal-$\uparrow$ Top-2 normal-$\uparrow$ Top-3 normal-$\uparrow$ FID normal-$\downarrow$ MM-DIST normal-$\downarrow$ Diversity normal-$\uparrow$

### v2

# CoMo: Controllable Motion Generation through Language Guided Pose Code Editing

@@ Abstract
Text-to-motion models excel at efficient human motion generation, but existing approaches lack fine-grained controllability over the generation process. Consequently, modifying subtle postures within a motion or inserting new actions at specific moments remains a challenge, limiting the applicability of these methods in diverse scenarios. In light of these challenges, we introduce CoMo, a Controllable Motion generation model, adept at accurately generating and editing motions by leveraging the knowledge priors of large language models (LLMs). Specifically, CoMo decomposes motions into discrete and semantically meaningful pose codes, with each code encapsulating the semantics of a body part, representing elementary information such as “left knee slightly bent”. Given textual inputs, CoMo autoregressively generates sequences of pose codes, which are then decoded into 3D motions. Leveraging pose codes as interpretable representations, an LLM can directly intervene in motion editing by adjusting the pose codes according to editing instructions. Experiments demonstrate that CoMo achieves competitive performance in motion generation compared to state-of-the-art models while, in human studies, CoMo substantially surpasses previous work in motion editing abilities. Project page: https://yh2371.github.io/como/.
@@ 1 Introduction
To address these challenges, we propose the Controllable Motion Generation model (CoMo). As shown in Figure 1, CoMo can generate high-quality human motions from a broad range of text (e.g., “A person throws a punch.”). Users can control the generation by altering the descriptions for each body part, e.g., switching the left and right arm descriptions changes the punching hand. More importantly, CoMo enables detailed editing across frames and joints (e.g., “Hunch over more.”), adding actions (e.g., “Squat at the end.”), and varying speed and even emotion (e.g., “more dramatic”), as depicted in Figure 6.
CoMo achieves these capabilities by representing motions as interpretable “pose codes”, with each code defining the state of a specific body part at a given moment, e.g., “right arm straight”. As demonstrated in Figure 2, our method starts by factorizing a motion sequence into a series of temporal states, each composed of pose codes that describe the motion’s kinematic characteristics. This structure allows for precise encoding of human motion sequences in time and fine-grained control of kinematic joint states. Since these encodings are explicitly interpretable, people can interact with and modify the sequences intuitively. We show that this process can also be made instructable with natural language by allowing large language models (LLM) to edit the codes as well.
Leveraging pose codes as interpretable motion representations, the three main components of CoMo work jointly to effectively generate and edit motion: (1) The Motion Encoder-Decoder applies heuristic rules to parse motions into sequences of semantically meaningful pose codes and trains a decoder to reconstruct these codes back into motions; (2) The Motion Generator, a transformer-based model, generates pose codes conditioned on text inputs and LLM-generated fine-grained descriptions; (3) The Motion Editor uses LLMs to modify and refine pose code sequences based on editing instructions. The resulting pose code sequences, whether generated or edited, are subsequently decoded into motion sequences using the previously trained decoder. CoMo allows for intuitive, language-controlled adjustments to the motion sequences, both temporally and kinematically, closely aligning the generated motions with users’ creative intentions and the nuances expressed in their textual descriptions, making the process user-friendly and adaptable to diverse applications.
@@ 2 Related Work
Fine-grained motion generation has gained significant interest due to its extensive practical applications. However, current methods have not fully integrated spatial and temporal details. TEACH [1] demonstrates improvements in motion smoothness by incorporating extensive temporal annotations but requires additional labeled data and overlooks lower-level spatial details. [17, 30, 2] leverage LLMs to generate detailed text descriptions for whole-body motions and individual body parts, aiming to map text to spatiotemporal details. However, without explicit supervision, these fine-grained details may not align with the respective motion. GraphMotion [16] proposes a hierarchical semantic graph that enforces coarse-to-fine topology within text-to-motion diffusion. However, constructing the hierarchical graph depends on the semantic parsing of text details and thus may struggle for ambiguous inputs. Also, the graph nodes do not model the lower-level spatial characteristics and relations. To this end, we propose body-part-specific semantic pose codes for achieving fine-grained representation of motion sequences. In addition, these pose codes are used as context to guide LLMs in generating motion-coherent descriptions for different body parts to enhance the connection between text and fine-grained motion details.
Motion Editing enables users to interactively refine generated motions to suit their expectations. PoseFix [6] automates 3D pose and text modifier generation for supervised editing, but frame-wise pose edits lack efficiency and temporal consistency. TLControl [35] allows motion editing using joint-level trajectories but relies on high-quality trajectory inputs and is not intuitive for interactive motion editing. [13] employs an autoencoder to optimize a motion manifold under positional, bone length, and trajectory constraints for smooth edits. Diffusion-based approaches [18, 9, 37] can achieve zero-shot spatiotemporal editing by infilling particular joints or frames, which may form unnatural discontinuities. Recently, FineMoGen [43] optimizes global attention for fine-grained editing but generates new sequences for each edit, limiting motion consistency. In our approach, we encode motions into semantic pose codes, which serve as context to prompt an LLM to edit the original motion directly, encouraging motion consistency.
@@ 3 Method
Figure 2: Overview of CoMo for text-driven motion generation. Motion Encoder-Decoder (left) utilizes a predefined codebook to encode motions into pose codes and learns a decoder to reconstruct the motions. Motion Generator (right), a transformer-based model, predicts pose codes autoregressively, conditioned on the text descriptions and LLM-generated fine-grained keywords. The generated pose codes are then decoded back into motions using the previously trained decoder.
@@ 3.1 Motion Encoder-Decoder
As shown in the left panel of Figure 2, given $T$ frames of motion $X=\left\{x_{i}\right\}_{i=1}^{T}$, the motion encoder $\mathcal{E}$ projects them into a sequence of pose codes using a codebook, denoted as $Z=\mathcal{E}\left(X\right)$, where $Z=\left\{z_{i}\right\}_{i=1}^{L}$. Here, $L=T/l$ represents the length of the pose code sequence, and $l$ is the downsampling rate over poses. The decoder $\mathcal{D}$ then decodes the pose codes back into motions, expressed as $X_{\text{rec}}=\mathcal{D}(\hat{Z})$, where $\hat{Z}$ stands for the latent features of pose codes. In the following, we will explain how to build the pose codebook, encode the motions using these codes, and learn the decoder to recover the motions.
Pose Codebook. Unlike [40], which uses an autoencoder to obtain implicit motion representations, CoMo predefines a semantic pose codebook to achieve interpretable motion factorization across time and space. Following PoseScript [5], we construct a pose codebook with $N$ codes, $\mathcal{C}=\left\{c_{n}\right\}_{n=1}^{N}$, with $c_{n}\in\mathbb{R}^{d_{c}}$, where $d_{c}$ is the dimension of the learnable codebook entry. Each code $c_{n}$ is associated with a semantic meaning, representing a state of a body part or spatial relationships between body parts, e.g., “left knee slightly bent”, “left hand and left foot close”. These codes are further grouped into $K$ pose categories, each encompassing different states of the same body parts, e.g., the pose category “left knee angle” includes codes like “left knee slightly bent”, “left knee partially bent”, etc.
Motion Encoder. Given the codebook $\mathcal{C}$ with $N$ pose codes grouped into $K$ categories, we encode motions into $K$-hot $N$ dimensional vectors, denoted as $Z=\mathcal{E}\left(X\right)=\left\{z_{i}\right\}_{i=1}^{L}$, where $z_{i}\in\mathbb{R}^{N}$. Here, K-hot indicates that $K$ elements of $z_{i}$ are set to $1$, while the others are 00. Specifically, at each time step $i$, we enforce mutual exclusivity between codes within the same category so that only one pose code per pose category is activated (set to $1$ in the vector). To determine whether a pose code $c\in\mathcal{C}$ applies to a motion $x$, we use an off-the-shelf skeleton parser [5], denoted as $\mathcal{P}(c,x)\rightarrow\{0,1\}$. The parser $\mathcal{P}$ analyzes the 3D joint positions of a skeleton in SMPL format [20], evaluating whether the pose meets specific heuristic threshold conditions for a given pose code. For instance, if the angle formed by the left shoulder, elbow, and wrist joints is less than 20 degrees, the code “left arm completely bent” is true. Therefore, instead of training an encoder model, we can explicitly factorize the motion sequence as:
@@
We define $V(X)=\left\{x_{i+1}-x_{i}\right\}_{i=1}^{T-1}$ as the velocity of the $T$-frame motion sequence $X$. The reconstruction objective is formulated with smooth L1 loss $\mathcal{L}_{1}$:
@@
where the hyperparameter $\lambda$ [^1] Following [40], we set $\lambda$ to 0.5. balances the velocity and motion loss. The learned codebook and decoder are then frozen for motion generation and editing.
@@ 3.2 Motion Generator
As illustrated in the right panel of Figure 2, the Motion Generator, conditioned on text input, aims to generate a sequence of pose codes, which will be decoded into motions. Utilizing the learned pose codebook, we map a motion sequence $X=\{x_{i}\}_{i=1}^{T}$ to a sequence of $K$-hot, $N$-dimensional vectors $Z^{1:L}=\left\{\{z_{i}^{n}\}_{n=1}^{N}\right\}_{i=1}^{L}$, where $z_{i}^{n}$ is an indicator function that is activated if the corresponding pose code $c_{n}$ is true at the time index $i$. To denote the end of a motion sequence, we append an <End> code to each latent vector, activated when the motion stops, and the dimension of each latent vector $z_{i}$ becomes $N+1$.
@@
To achieve this goal, we adopt a decoder-only transformer architecture with causal self-attention [40]. [^2] Following the method for processing image patches in Vision Transformers [7], a linear layer projects the $K$-hot vectors before input into the transformer architecture. The likelihood of the full sequence is:
@@
The predicted sequences of $K$-hot vectors are mapped back to respective pose codes, which can then be decoded to motion sequences through the decoder $\mathcal{D}$.
Fine-grained Keywords. To help the model capture more fine-grained details, we enhance the text description by using GPT-4 [24] to generate a keyword for each of the 10 body parts [^3] The 10 body parts are head, torso, left arm, right arm, left hand, right hand, left leg, right leg, left feet, right feet. and a keyword to describe the overall mood of the motion. For example, as shown in Figure 2, GPT-4 generates “bend at the knee”, “torso leaning forward”, etc., as the keywords for the motion “crouching”. We use CLIP [27] to extract text embeddings of the keywords and original description. The embedding of the original text, followed by the embeddings of 11 keywords, form the initial token sequence that conditions the subsequent motion generation. The effectiveness of using these keywords to improve generation performance is validated in Table 3.
@@ 3.3 Motion Editor
The edited segments of the pose codes are seamlessly integrated with their unedited counterparts, forming the complete edited sequence of pose codes, which is passed through the decoder $\mathcal{D}$ to reconstruct the final edited motion.
@@ 4 Evaluation
In this section, we evaluate CoMo from two perspectives: (1) experiments to demonstrate that CoMo achieves state-of-the-art performance on text-driven motion generation (Sec 4.1), and (2) human evaluation to showcase that CoMo is superior over existing methods in motion editing (Sec 4.2). We also conduct comprehensive ablation studies (Sec 4.3) to validate our model design.
@@ 4.1 Experiments on Motion Generation
Method R-Precision $\uparrow$ FID $\downarrow$ MM-DIST $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
@@
Table 2: Comparison with the state-of-the-art methods on KIT [26] test set. The best performance is bold, and the second best is underlined, the third best is italic.
Method R-Precision $\uparrow$ FID $\downarrow$ MM-DIST $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
@@
CoMo achieves competitive results on text-driven motion generation. Table 1 and Table 2 present the results of text-driven motion generation on the HumanML3D and KIT-ML datasets, respectively. As shown in the first two rows of both tables, motion reconstruction with discrete pose codes achieves high fidelity and closely matches ground-truth text-motion consistency, providing a strong foundation for motion generation. CoMo attains either the best or the second-best performance on HumanML3D across five metrics and ranks within the top three in six metrics on KIT. Compared to existing state-of-the-art methods, CoMo not only achieves competitive motion fidelity (FID) but also enhances motion diversity and consistency between generated motion and text descriptions (R-Precision, MM-DIST, Diversity). MModality measures the diversity of motions generated from the same text description. Due to the use of semantically meaningful pose codes, CoMo builds a stronger binding between text and the generated motion sequence, which potentially causes a slightly lower MModality score compared to prior methods as a trade-off for consistency. More importantly, our approach provides an intuitive interface for LLM-based zero-shot motion editing, which will be evaluated in Section 4.2 via a user study.
@@ 4.2 Human Evaluation on Motion Editing
Baselines. To form strong baselines, we prompt GPT-4 to generate an updated description in the style of HumanML3D annotations using the original description and edit instructions as context [^7] Prompts and updated descriptions are available in the Appendix.. Wcription to generate the edited motion for T2M-GPT and FineMoGen. For CoMo, we decompose the source motion as pose codes and prompt GPT-4 to edit the code sequence based on the original description, edit instruction and pose code semantics without fine-tuning. The edited code sequence is decoded to reconstruct the edited motion.
@@
Figure 6: Qualitative examples of Motion Editing on the HumanML3D test set [10]. The green words/boxes highlight successful edits. The red words/boxes identify misalignments between edited and source motions. Compared to other methods, CoMo achieves accurate edits while preserving key characteristics of the source motion.
@@
Top-1 $\uparrow$ FID $\downarrow$ MM-DIST $\downarrow$ Diversity $\uparrow$ Top-1 $\uparrow$ FID $\downarrow$ MM-DIST $\downarrow$ Diversity $\uparrow$
@@
Top-1 $\uparrow$ FID $\downarrow$ MM-DIST $\downarrow$ Diversity $\uparrow$
@@
Sampling Rate ($l$) Top-1 $\uparrow$ Top-2 $\uparrow$ Top-3 $\uparrow$ FID $\downarrow$ MM-DIST $\downarrow$ Diversity $\uparrow$
@@
Humans prefer CoMo for Motion Editing. We present the average percentage of users’ preferences for CoMo versus T2M-GPT and FineMoGen in Figure 5. The results show a clear preference (over 70% on average) for motion editing with CoMo, especially in scenarios thatdify fine-grained motion details (e.g. body part modification and add/delete action). As depicted in Figure 6, while T2M-GPT and FineMoGen can produce motion relevant to the updated description, they often struggle to generate well-aligned details from scratch without any spatial-temporal context of the source motion. In contrast, CoMo leverages the semantics of pose codes to interpret the source motion effectively and achieve detailed motion editing. The observed decrease in preference for holistic motion edits involving changes in emotion or speed suggests that textual descriptions may provide more guidance in creating comprehensive edits that impact multiple aspects of a motion sequence. With its capability for high-level text-guided generation and iterative fine-grained pose code editing, CoMo emerges as a competitive approach for text-driven, controllable motion generation.
4.3 Ablation
Codebook Size. The number of pose codes indicates how fine-grained the heuristic thresholds of the skeleton parser are. More fine-grained thresholds capture more details at the expense of computational speed and learning complexity. We evaluate motion reconstruction quality for different codebook sizes by varying the angle and distance threshold settings, with finer thresholds corresponding to larger codebooks. Results in Table 4 indicate that our current codebook setting of $392$ pose codes best balances complexity and reconstruction quality.